# Measuring the Speed of Gravity and the Cosmic Expansion with Time Delays between Gravity and Light from Binary Neutron Stars

### v1

@@ Abstract
The first observation of a gravitational wave (GW) and a short gamma-ray burst (sGRB) emitted by the same binary neutron star (BNS) merger officially opened the field of GW multi-messenger astronomy. In this paper, we define and address lagging sirens, a new class of multi-messenger BNSs for which associated GWs and sGRBs are observed without the identification of their host galaxy. We propose a new methodology to use the observed time delay of these sources to constrain the speed of gravity that is, the propagation speed of gravitational waves, the Hubble constant and the prompt time delay distribution between GWs and sGRBs, even though a direct redshift estimation from the host galaxy is unavailable. Our method exploits the intrinsic relation between GWs and sGRBs observed and prompt time delays to obtain a statistical redshift measure for the cosmological sources. We show that this technique can be used to infer the Hubble constant at the $10\%$ level of precision with future-generation GW detectors such as the Einstein Telescope and only 100 observations of this kind. The novel procedure that we propose has systematics that differ completely from the ones of previous GW methods for cosmology. Additionally, we demonstrate for the first time that the speed of gravity and the distribution of the prompt time-delays between GWs and sGRBs can be inferred conjointly with less than 10 sources even with current GW detector sensitivities.
@@ I Introduction
So far, GW170817 is the only GW event with an unambiguously associated EM counterpart. Out of almost 100 GW detections achieved to date [9, 10, 11], most are due to binary black hole mergers observed without an EM counterpart (dark sirens). Among the several techniques proposed for dark sirens cosmology is the source-frame mass method [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. This method consists of exploiting the difference between the detector mass $m_{d}$ and source mass $m_{s}$ of CBCs, which are related by a redshift factor $m_{d}=(1+z)m_{s}$. Thus, by knowing the source-frame mass spectrum, it is possible to infer statistically the source redshift. Since the source mass spectrum is not known accurately, current studies conjointly fit flexible phenomenological models of the mass spectrum along with cosmological parameters that can in some limits approximate the true binary distributions [25]. The source frame mass method was used with the latest GW detections to obtain a value of $H_{0}=68^{+12}_{-7}{\rm km\,s^{-1}\,Mpc^{-1}}$ [26]. The low precision in the inferred value of $H_{0}$ is due the broadness of the source-frame mass spectrum, which prevents us from obtaining precise redshift estimations for the GW event sources used to measure $H_{0}$.
@@
The paper is organized as follows. In Sec. II, we introduce the relations between prompt and observed GW-GRB time delays considering also possible deviations between the speed of light and gravity. We further introduce the statistical framework used to infer the Hubble constant $H_{0}$, speed of gravity $\hat{\alpha}_{0}$ and prompt time delay distribution. In Sec. III simulate catalogs of GW and sGRB observations considering two scenarios for the sensitivity of GW detectors, namely, the sensitivity expected to be achieved in 2027-2030 with the fifth LIGO-Virgo-KAGRA Observing Run (O5) [31, 32, 33, 34] and an ET-like sensitivity expected to be achieved in 2035+. In Sec. IV we present the prospects of measuring $H_{0}$, $\hat{\alpha}_{0}$ and the prompt time delay distribution. Finally, in Sec. V we draw our conclusions.
@@ II.1 Theoretical background
We define the prompt time delay $\Delta t_{\rm s}$ as the elapsed time between the GW luminosity peak at the CBC merger and the emission of the sGRB. A positive $\Delta t_{\rm s}$ indicates that the sGRB is emitted after the CBC merger, while a negative value indicates that is emitted prior to the merger. The observed time delay between GW and GRB at the merger is [35]
@@
In writing Eq. (2), we implicitly assumed that the speed of gravity does not depend on the GW frequency and that its deviation from the speed of light is small [35]. Moreover, we assumed that the speed of gravity does not change during cosmic time, although this possibility is not excluded in principle for some theories of gravity [36, 37].
@@
Equation (2) can be used to predict the observed time delays between the GW and sGRB signals from the same event. For instance, if the GW and sGRB were prompted together ($\Delta t_{\rm s}=0$ s), and GWs were slightly slower than the light ($\hat{\alpha}_{0}<0$), then the GW would be detected after the arrival of the sGRB ($\Delta t_{\rm d}<0$ s). If instead the sGRB were prompted after the GW ($\Delta t_{\rm s}>0$ s) and the speed of gravity were equal to $c$ ($\hat{\alpha}_{0}=0$), then the GW would surely be detected prior to the sGRB ($\Delta t_{\rm d}>0$ s). Figure 1 shows the observed time delays as a function of redshift for various values of the speed of gravity and prompt time delay. The larger the redshift, the longer the observed time delay between the GW and the sGRB is. From Fig. 1, we can already see that if the observed time delay were accurately measured [^1] This is a reasonable assumption also supported by the measure of GW170817 with its sGRBs for which an observed time delay of $\Delta t_{\rm d}=1.74\pm 0.05$ s was observed [4]. and the prompt time delay and speed of gravity were known, one could get implicit redshift information for the GW/sGRB sources. Unfortunately, while we can argue that the speed of gravity must be equal to the speed of light if we assume classical General Relativity, the distribution of prompt time delays between GWs and sGRBs is not a constant and it is currently unknown. For this reason, we introduce a statistical framework able to infer it.
@@ II.2 Statistical background
Let us assume that we have a set of common population parameters $\vec{\lambda}$, such as the Hubble constant or the propagation speed of GWs, that we want to infer from $N_{\rm obs}$ GW+sGRB joint observations collected from the data $\vec{d}_{1},\ldots,\vec{d}_{N_{\rm obs}}$}. We further assume that each astrophysical source is described by a set of parameters $\vec{\theta}$ the distributions of which are described by a CBC rate model
@@
where $t_{\rm d}$ is the detector time. Our aim is to infer the parameters $\vec{\lambda}$ by conjointly fitting the rate distribution of the parameters $\vec{\theta}$. The hierarchical likelihood of obtaining these detections is [38]:
@@
For this study, the relevant $\vec{\theta}$ parameters are the luminosity distance $D_{\rm L}$ and the observed GW-sGRB time delay $\Delta t_{\rm d}$. Thus, the CBC rate model is written in terms of these two quantities is
@@
$\frac{\partial D_{\rm L}}{\partial z}=\frac{D_{\rm L}(z)}{1+z}+\frac{c(1+z)}{H_{0}}\frac{1}{\sqrt{\Omega_{m}(1+z^{\prime})^{3}+(1-\Omega_{m})}}\,.$ (10)
@@
where $R_{0}$ is the BNS merger rate per comoving volume at redshift $z=0$, $\psi(z;\vec{\lambda})$ is a phenomenological parametrization for the BNS merger rate as a function of redshift such that $\psi(z=0;\Lambda)=1$, $p_{\rm pop}$ is a probability density function describing the prompt time delay distribution, and
@@ II.3 Study case with GW170817 and GRB170817A
where $\mathcal{L}(\vec{d}_{i}^{\rm EM}|z,\sigma_{z})$ encapsulates the measure of the redshift from the source host galaxy. Following [6], we assume that this is a Gaussian distribution with mean $z=0.0100$ and $\sigma=0.0005$. Note also that in Eq. (15), the detection probability at the denominator should also contain the detection probability of the EM counterpart. However, we neglect this component as during the O2 run, the detection range of electromagnetic counterparts was significantly higher than the one of the GW sources.
To estimate the source luminosity distance, and hence the redshift, we use the low-spin parameter estimation samples released [^2] https://dcc.ligo.org/LIGO-P1800061/public with [7]. The estimated value of the luminosity distance from the posterior samples is $D_{\rm L}=41^{+6}_{-12}$ Mpc [7]. We use a flat $\Lambda$CDM model with cosmological parameters from [40] ($H_{0}=67.66\,{\rm km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{m}=0.310$) to convert between luminosity distance and redshift. We also assume that the time delay at the detector between GW170817 and GRB170817 is $\Delta t_{d}=1.74^{+0.05}_{-0.05}$ s, measured with a Gaussian posterior. More importantly, we fix the distribution of source prompt time delays to be uniform between [0,10] s, meaning that the GRB can be emitted up to 10 seconds after the merger. This choice is made following [4], where to derive the constraints on the speed of gravity it was assumed that the GRB could be emitted at most 10 seconds after the merger.
Figure 2: Posterior distributions for the speed of gravity from the joint observation of GW170817-GRB170817A. The green line represents the posterior distribution using only the measurement of Luminosity Distance ($D_{\rm L}$) and Time Delay ($\Delta t_{\rm d}$). The blue line represents the posterior distribution when including the redshift information from the event host galaxy.
Figure 2 shows the posterior distribution of the fractional difference between the speed of gravity and the speed of light. The boundaries of the posterior can be used as a proxy for the constraints on the speed of light. When just using GW170817 and GRB170817A, we obtain
@@
which match the constraints obtained in [4]. When we also include the information on the redshift of NG4993, we obtain
@@
We can see that the posterior generated including the redshift estimation of NGC4993 has shorter tails than the one generated without redshift. This is due to the fact that, the redshift of GW170817 is more precise the the luminosity distance inferred from the GW. As a consequence, when we fix the cosmology, the luminosity distance of the source is almost entirely given by the galaxy host redshift, with a negligible uncertainty on its value.
@@ III Mock Catalogs of GW and sGRB detections
We simulated mock catalogues of GW and sGRB detections to forecast the potential of the method we propose to constrain prompt time delay distributions, speed of gravity and $H_{0}$. In Sec. III.1 we describe how we simulate joint detections of GW and sGRB, while in Sec. III.2 we provide a general overview of the observational properties in terms of luminosity distance and observed time delay of the GW/sGRB detections.
@@ III.1 Catalogs generation
In Fig. 3 we depict the simulation procedure used to generate the GW/sGRB mock catalogs.
@@
We choose a flat $\Lambda\text{CDM}$ cosmology with parameters matching those derived from the latest observations of the Cosmic Microwave Background by Planck [40] ($H_{0}=67.66\,{\rm km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{m}=0.310$) and we also assume that GWs propagate at the speed of light ($\hat{\alpha}_{0}=0$). We simulate BNS sources with an isotropic distribution over the sky in terms of right ascension $\alpha$ and declination $\delta$. We also simulate an isotropic distribution for the binary orbital inclination angle $\iota$ and a uniform distribution of polarization angles $\psi$. The source masses $m_{1}$ and $m_{2}$ of the binary neutron stars are drawn uniformly in the interval $(1,3)\,M_{\odot}$, with the condition $m_{2}<m_{1}$. This distribution is consistent with the latest rate models inferred from the GW detections during the third LIGO-Virgo-KAGRA Observing Run (O3) [41]. The spin magnitudes of the two neutron stars are assumed to be vanishing. This is consistent with expectations for BNS, which are low-spin objects [7]. The distribution of the BNS sources in redshift is modelled according to a Madau-Dickinson-like star formation rate [42], namely
@@
In this work, we model the GW SNR $\rho^{2}_{\text{opt},i}$ at the detector $i$ at the 0-th Post Newtonian order [44], namely
@@
with $\mathcal{M}_{\text{obs}}$ the redshifted chirp mass. $F_{+,i}$ and $F_{\times,i}$ are the detector antenna patterns for the detector $i$, for a GW signal located at right ascension $\alpha$ and declination $\delta$, with polarization angle $\psi$, and arriving at a time $t$. The quantity $I_{i}$ is defined as:
@@
with $\theta_{B}=4.5\,$rad, compatible with the average value of the opening angle as found by [46]), and $L_{\text{max}}$ is drawn from a log-normal distribution with a mean of $5\times 10^{51}$ erg s-1 and width of 0.56 dex [45]. The simple parametrization in Eq. (23) predicts a small luminosity if the binary is observed edge-on $\iota=\pi/2\gg\theta_{B}$ and a luminosity equal to $L_{\text{max}}$ when the binary is observed face-on ($\iota=0\,$rad). This model is consistent with most of the observed sGRBs with known redshift and jet opening angle [47, 4]. To compute the sGRB peak energy in the source frame $E_{p}$, we model a relation between $E_{p}$ and $L(\iota;\theta_{B},L_{\text{max}})$ in analogy to the one of long bursts [48]. Following [46]; we write this correlation in the form:
@@
To represent measurement uncertainties inherent in actual observations, we generate simulated error values for the sources’ luminosity distances and observed GW/sGRB time delays. The measurement of the luminosity distance is provided from the GW detection. In principle, to obtain reliable distance estimates, one should use a full Bayesian analysis using simulated data and implementing state-of-the-art waveform models for the GW signals. However, this approach is highly computationally demanding. For this reason, we adopt a GW likelihood [51] model able to describe statistically the typical error budgets for the GW luminosity distance. With this model, we associate a set of data $x_{i}$ with an “observed luminosity distance” $D_{L}^{\text{obs}}$, namely
@@
where $D_{L}^{\text{true}}$ is the true physical value of the luminosity distance. Additionally, we also assume that the observed SNR is not the true SNR in Eq. (20) but it is modified by possible noise fluctiations. To mimic this effect, we generate the observed SNR $\rho_{\text{obs}}$ from a Gaussian distribution with a mean equal to the intrinsic SNR $\rho_{\text{true}}$ and a standard deviation of 1 [52] and apply the SNR selection criterion on it. The likelihood for the observed SNR is

### v2

@@ Abstract
The first observation of a gravitational wave (GW) and a short gamma-ray burst (sGRB) emitted by the same binary neutron star (BNS) merger officially opened the field of GW multimessenger astronomy. In this paper, we define and address lagging sirens, a new class of multimessenger BNSs for which associated GWs and sGRBs are observed without the identification of their host galaxy. We propose a new methodology to use the observed time delay of these sources to constrain the speed of gravity that is, the propagation speed of gravitational waves, the Hubble constant and the prompt time delay distribution between GWs and sGRBs, even though a direct redshift estimation from the host galaxy is unavailable. Our method exploits the intrinsic relation between GWs and sGRBs observed and prompt time delays to obtain a statistical redshift measure for the cosmological sources. We show that this technique can be used to infer the Hubble constant at the $10\%$ level of precision with future-generation GW detectors such as the Einstein Telescope and only 100 observations of this kind. The novel procedure that we propose has systematics that differ completely from the ones of previous GW methods for cosmology. Additionally, we demonstrate for the first time that the speed of gravity and the distribution of the prompt time delays between GWs and sGRBs can be inferred conjointly with less than 10 sources even with current GW detector sensitivities.
@@ I Introduction
So far, GW170817 is the only GW event with an unambiguously associated EM counterpart. Out of almost 100 GW detections achieved to date [9, 10, 11], most are due to binary black hole mergers observed without an EM counterpart (dark sirens). Among the several techniques proposed for dark sirens cosmology is the source-frame mass method [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. This method consists of exploiting the difference between the detector mass $m_{d}$ and source mass $m_{s}$ of CBCs, which are related by a redshift factor $m_{d}=(1+z)m_{s}$. Thus, by knowing the source-frame mass spectrum, it is possible to infer statistically the source redshift. Since the source mass spectrum is not known accurately, current studies conjointly fit flexible phenomenological models of the mass spectrum along with cosmological parameters that can in some limits approximate the true binary distributions [25]. The source frame mass method was used with the latest GW detections to obtain a value of $H_{0}=68^{+12}_{-7}{\rm km\,s^{-1}\,Mpc^{-1}}$ [26]. The low precision in the inferred value of $H_{0}$ is due to the broadness of the source-frame mass spectrum, which prevents us from obtaining precise redshift estimations for the GW event sources used to measure $H_{0}$.
@@
The paper is organized as follows. In Sec. II, we introduce the relations between prompt and observed GW-sGRB time delays considering also possible deviations between the speed of light and gravity. We further introduce the statistical framework used to infer the Hubble constant $H_{0}$, speed of gravity $\hat{\alpha}_{0}$ and prompt time delay distribution. In Sec. III we simulate catalogs of GW and sGRB observations considering two scenarios for the sensitivity of GW detectors, namely, the sensitivity expected to be achieved in 2027-2030 with the fifth LIGO-Virgo-KAGRA Observing Run (O5) [31, 32, 33, 34] and an ET-like sensitivity expected to be achieved in 2035+. In Sec. IV we present the prospects of measuring $H_{0}$, $\hat{\alpha}_{0}$ and the prompt time delay distribution. Finally, in Sec. V we draw our conclusions.
@@ II.1 Theoretical background
We define the prompt time delay $\Delta t_{\rm s}$ as the elapsed time between the GW luminosity peak at the CBC merger and the emission of the sGRB. A positive $\Delta t_{\rm s}$ indicates that the sGRB is emitted after the CBC merger, while a negative value indicates that it is emitted prior to the merger. The observed time delay between GW and sGRB at the merger is [35]
@@
In writing Eq. (2), we implicitly assumed that the speed of gravity does not depend on the GW frequency and that its deviation from the speed of light is small [35]. Moreover, we assumed that the speed of gravity does not change during cosmic time, although this possibility is not excluded in principle for some theories of gravity [36, 37] or agnostic tests on the speed of gravity [38].
@@
Equation (2) can be used to predict the observed time delays between the GW and sGRB signals from the same event. For instance, if the GW and sGRB were prompted together ($\Delta t_{\rm s}=0$ s), and GWs were slightly slower than the light ($\hat{\alpha}_{0}<0$), then the GW would be detected after the arrival of the sGRB ($\Delta t_{\rm d}<0$ s). If instead the sGRB were prompted after the GW ($\Delta t_{\rm s}>0$ s) and the speed of gravity were equal to $c$ ($\hat{\alpha}_{0}=0$), then the GW would surely be detected prior to the sGRB ($\Delta t_{\rm d}>0$ s). Figure 1 shows the observed time delays as a function of redshift for various values of the speed of gravity and prompt time delay. The larger the redshift, the longer the observed time delay between the GW and the sGRB is. From Fig. 1, we can already see that if the observed time delay were accurately measured [^1] This is a reasonable assumption also supported by the measure of GW170817 with its sGRBs for which an observed time delay of $\Delta t_{\rm d}=1.74\pm 0.05$ s was observed [4]. and the prompt time delay and speed of gravity were known, one could get implicit redshift information for the GW-sGRB sources. Unfortunately, while we can argue that the speed of gravity must be equal to the speed of light if we assume classical general relativity, the distribution of prompt time delays between GWs and sGRBs is not a constant and it is currently unknown. For this reason, we introduce a statistical framework able to infer it.
@@ II.2 Statistical background
Let us assume that we have a set of common population parameters $\vec{\lambda}$, such as the Hubble constant or the propagation speed of GWs, that we want to infer from $N_{\rm obs}$ GW-sGRB joint observations collected from the data $\vec{d}_{1},\ldots,\vec{d}_{N_{\rm obs}}$}. We further assume that each astrophysical source is described by a set of parameters $\vec{\theta}$, the distributions of which are described by a CBC rate model
@@
where $t_{\rm d}$ is the detector time. Our aim is to infer the parameters $\vec{\lambda}$ by conjointly fitting the rate distribution of the parameters $\vec{\theta}$. The hierarchical likelihood of obtaining these detections is [39]
@@
For this study, the relevant $\vec{\theta}$ parameters are the luminosity distance $D_{\rm L}$ and the observed GW-sGRB time delay $\Delta t_{\rm d}$. Thus, the CBC rate model written in terms of these two quantities is
@@
$\frac{\partial D_{\rm L}}{\partial z}=\frac{D_{\rm L}(z)}{1+z}+\frac{c(1+z)}{H_{0}}\frac{1}{\sqrt{\Omega_{m}(1+z)^{3}+(1-\Omega_{m})}}\,.$ (10)
@@
where $R_{0}$ is the BNS merger rate per comoving volume at redshift $z=0$, $\psi(z;\vec{\lambda})$ is a phenomenological parametrization for the BNS merger rate as a function of redshift such that $\psi(z=0;\vec{\lambda})=1$, $p_{\rm pop}$ is a probability density function describing the prompt time delay distribution, and
@@ II.3 Study case with GW170817 and GRB 170817A
where $\mathcal{L}(\vec{d}_{i}^{\rm EM}|z,\sigma_{z})$ encapsulates the measure of the redshift from the source host galaxy. Following [6], we assume this is a Gaussian distribution with mean $z=0.0100$ and $\sigma=0.0005$. Note also that in Eq. (15), the detection probability at the denominator should also contain the detection probability of the EM counterpart. However, we neglect this component as during the O2 run, the detection range of electromagnetic counterparts was significantly higher than the one of the GW sources. To estimate the source luminosity distance, and hence the redshift, we use the low-spin parameter estimation samples released [^2] https://dcc.ligo.org/LIGO-P1800061/public with [7]. The estimated value of the luminosity distance from the posterior samples is $D_{\rm L}=41^{+6}_{-12}$ Mpc [7]. We use a flat $\Lambda$CDM model with cosmological parameters from [41] ($H_{0}=67.66\,{\rm km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{m}=0.310$) to convert between luminosity distance and redshift. We also assume that the time delay at the detector between GW170817 and GRB 170817A is $\Delta t_{d}=1.74^{+0.05}_{-0.05}$ s, measured with a Gaussian posterior. More importantly, we fix the distribution of source prompt time delays to be uniform between [0,10] s, meaning that the GRB can be emitted up to 10 seconds after the merger. This choice is made following [4], where to derive the constraints on the speed of gravity it was assumed that the GRB could be emitted at most 10 seconds after the merger.
Figure 2: Posterior distributions for the speed of gravity from the joint observation of GW170817 and GRB 170817A. The green line represents the posterior distribution using only the measurement of luminosity Distance ($D_{\rm L}$) and Time Delay ($\Delta t_{\rm d}$). The blue line represents the posterior distribution when including the redshift information from the event host galaxy.
Figure 2 shows the posterior distribution of the fractional difference between the speed of gravity and the speed of light. The boundaries of the posterior can be constrained as a proxy for the constraints on the speed of gravity. When just using GW170817 and GRB 170817A, we obtain
@@
which matches the constraints obtained in [4] by combining only GW and Fermi-INTEGRAL data. Notably, in [4], the authors excluded inputs from external analyses, such as the identification of NGC 4993 via EM observations at lower wavelengths, and avoided assumptions about cosmology required to convert the luminosity distance into redshift. Instead, they assumed a conservative (in terms of speed of gravity constraints) lower limit of 26 Mpc for the luminosity distance. When we also include the information on the redshift of NGC 4993, we obtain
@@
We can see that the posterior generated including the redshift estimation of NGC 4993 has shorter tails than the one generated without redshift. This is due to the fact that, the redshift of GW170817 is more precise than the luminosity distance inferred from the GW. As a consequence, when we fix the cosmology, the luminosity distance of the source is almost entirely given by the galaxy host redshift, with a negligible uncertainty on its value.
@@ III Mock Catalogs of GW and sGRB detections
We simulated mock catalogs of GW and sGRB detections to forecast the potential of the method we propose to constrain prompt time delay distributions, speed of gravity and $H_{0}$. In Sec. III.1 we describe how we simulate joint detections of GW and sGRB, while in Sec. III.2 we provide a general overview of the observational properties in terms of luminosity distance and observed time delay of the GW-sGRB detections.
@@ III.1 Catalogs generation
In Fig. 3 we depict the simulation procedure used to generate the GW-sGRB mock catalogs.
@@
We choose a flat $\Lambda\text{CDM}$ cosmology with parameters matching those derived from the latest observations of the cosmic microwave background by Planck [41] ($H_{0}=67.66\,{\rm km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{m}=0.310$) and we also assume that GWs propagate at the speed of light ($\hat{\alpha}_{0}=0$). We simulate BNS sources with an isotropic distribution over the sky in terms of right ascension $\alpha$ and declination $\delta$. We also simulate an isotropic distribution for the binary orbital inclination angle $\iota$ and a uniform distribution of polarization angles $\psi$. The source masses $m_{1}$ and $m_{2}$ of the binary neutron stars are drawn uniformly in the interval $(1,3)\,M_{\odot}$, with the condition $m_{2}\leq m_{1}$. This distribution is consistent with the latest rate models inferred from the GW detections during the third LIGO-Virgo-KAGRA Observing Run (O3) [42]. The spin magnitudes of the two neutron stars are assumed to be vanishing. This is consistent with expectations for BNS, which are low-spin objects [7]. The distribution of the BNS sources in redshift is modeled according to a Madau-Dickinson-like star formation rate [43], namely
@@
In this work, we model the GW SNR $\rho_{\text{opt},i}$ at the detector $i$ at the 0th post-Newtonian order [45], namely
@@
with $\mathcal{M}_{\text{obs}}$ the redshifted chirp mass. $F_{+,i}$ and $F_{\times,i}$ are the detector antenna patterns for the detector $i$, for a GW signal located at right ascension $\alpha$ and declination $\delta$, with polarization angle $\psi$, and arrival at a time $t$. In our simulations, we fixed the antenna pattern functions at the arrival time $t$ of the GW. This approximation is reasonable for the current ground-based detector, as the BNS signals enter the sensitivity band around $10$ Hz, with only a few minutes left prior to the merger. Since the antenna patterns vary with the sidereal day, they can be considered constant for this case. However, for ET, BNS signals will enter the sensitivity band around $\sim O(1)$ Hz and be visible for a few days. In order to ease the computational load of our simulations, we impose a lower limit to the ET sensitivity at $10$ Hz. In other words, we are artificially reducing the GW detection range that ET would have. However, as we will see later, this is not a problem for the simulation as the detection range for GRBs with Fermi-like satellites is significantly smaller than the restricted ET detection range.
The quantity $I_{i}$ is defined as
@@
with $\theta_{B}=4.5\,$rad, compatible with the average value of the opening angle as found by [47], and $L_{\text{max}}$ is drawn from a log-normal distribution with a mean of $5\times 10^{51}$ erg s-1 and width of 0.56 dex [46]. The simple parametrization in Eq. (23) predicts a small luminosity if the binary is observed edge-on $\iota=\pi/2\gg\theta_{B}$ and a luminosity equal to $L_{\text{max}}$ when the binary is observed face-on ($\iota=0\,$rad). This model is consistent with most of the observed sGRBs with known redshift and jet opening angle [48, 4]. To compute the sGRB peak energy in the source frame $E_{p}$, we model a relation between $E_{p}$ and $L(\iota;\theta_{B},L_{\text{max}})$ in analogy to the one of long bursts [49]. Following [47], we write this correlation in the form
@@
To represent measurement uncertainties inherent in actual observations, we generate simulated error values for the sources’ luminosity distances and observed GW-sGRB time delays. The measurement of the luminosity distance is provided from the GW detection. In principle, to obtain reliable distance estimates, one should use a full Bayesian analysis using simulated data and implementing state-of-the-art waveform models for the GW signals. However, this approach is highly computationally demanding. For this reason, we adopt a GW likelihood [52] model able to describe statistically the typical error budgets for the GW luminosity distance. With this model, we associate a set of data $x_{i}$ with an “observed luminosity distance” $D_{L}^{\text{obs}}$, namely
@@
where $D_{L}^{\text{true}}$ is the true physical value of the luminosity distance. Additionally, we also assume that the observed SNR is not the true SNR in Eq. (20) but it is modified by possible noise fluctuations. To mimic this effect, we generate the observed SNR $\rho_{\text{obs}}$ from a Gaussian distribution with a mean equal to the intrinsic SNR $\rho_{\text{true}}$ and a standard deviation of 1 [53] and apply the SNR selection criterion on it. The likelihood for the observed SNR is